jellyfish are about 19 water. if you have a 300 lb jellyfish( and yes, some do
--
20
grow to be that big!) what is the weight of the water it contains?(show your work, but you do not need a 5 step solution.)

Answer by stanbon(75887)

jellyfish are about 19 water. if you have a 300 lb jellyfish( and yes, some do
--
20
grow to be that big!) what is the weight of the water it contains?(show your work, but you do not need a 5 step solution.)
------------------------------------------------
x/300 = 19/20
x = (19/20)*300
---
x = 19*15
x = 285 lbs of water
============================
Cheers,
Stan H.